Fdegrees=9/5 C+32 C=15degrees sorry can't get the little degree sign up where it should go. I don't know how to do this.I did 15 X 9/5 + 32 and I get 59 but my teacher says the answer is 48 4/9 What did I do wrong. I have another problem just like this one so your answer will help me with both
Found 2 solutions by sarahschober, stanbon:
Answer by sarahschober(35)
Answer by stanbon(75887)

Fdegrees=9/5 C+32 C=15degrees sorry can't get the little degree sign up where it should go. I don't know how to do this.I did 15 X 9/5 + 32 and I get 59
-------------
F=(9/5)C+32
F(15)=(9/5)(15)+32
F(15)=27+32
f(15)=59
I think you are right.
Cheers,
Stan H.